OpenStudy (anonymous):

Gobs of Globes is a specialty store which carries a variety of globe-related products. One of their popular products is a globe piggy bank. The standard one they carry has a volume of 500/3 cubic inches. A customer has asked for a custom-made version with a radius double the size of the standard one. Based on this information, what will be the radius of the custom-made version? 25??

OpenStudy (apoorvk):

The volume 'V' of a sphere of radius 'r' is: \[V=\frac{4}{3}\pi (r)^3\] Now what happens to the volume when 'r' becomes '2r'?

OpenStudy (apoorvk):

@suzy4321

OpenStudy (anonymous):

hm im confused i know radius is 6 x6 example but 2r

OpenStudy (apoorvk):

Okay, what is (2r)^3?

OpenStudy (anonymous):

8r

OpenStudy (apoorvk):

It's = 2r*2r*2r = 8r^3 ---> right? yeah, so 8r^3 is 8 times of r^3 - am i correct?

OpenStudy (apoorvk):

so, when I am doubling the radius, the volume becomes '8' times the initial one - do you understand how?

OpenStudy (anonymous):

yes like going backwads

OpenStudy (apoorvk):

'backwards'? you mean back-calculating? yeah, in a way.. so, what is the new volume then?

OpenStudy (apoorvk):

???

OpenStudy (anonymous):

4/3r3 500/3 in= 4/3r3 125 in3 = r3 5 in = r soo ten

OpenStudy (anonymous):

so much rushing for that jaja

OpenStudy (apoorvk):

oops lol, i was thinking all the time that we need to find the new volume :/ Yeah, you're right, @suzy4321